OpenStudy (anonymous):

how do you solve 3+6/t-3=6/t^2-3t?

OpenStudy (anonymous):

Can you make the factor of denominator of Right Hand Side??

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

would it be t(t-3) factored?

OpenStudy (anonymous):

Yes you are right.. There is one (t-3) on left hand side now can you cancel them??

OpenStudy (anonymous):

\[\frac{9}{t-3} = \frac{6}{t(t-3)} \implies \frac{9}{\cancel{(t-3)}} = \frac{6}{t \cancel{(t-3)}}\] Can you find t now??

OpenStudy (anonymous):

Wait what is your question??

OpenStudy (anonymous):

let me see...

OpenStudy (anonymous):

how do you solve 3+6/t-3=6/t^2-3t?

OpenStudy (anonymous):

Is it??